Mullineaux Corporation has a target capital structure of 75 percent common stock, 5 percent preferred stock, and 20 percent debt. Its cost of equity is 11.25 percent, the cost of preferred stock is 5.5 percent, and the cost of debt is 6.1 percent. The relevant tax rate is 35 percent.

Required:
(a)

What is Mullineaux’s WACC? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).)

  WACC %
(b)

What is the aftertax cost of debt? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).)

  Aftertax cost of debt %
Answer #1

After-tax cost of debt=6.1*(1-tax rate)

=6.1*(1-0.35)=3.965%

=3.97%(Approx).

WACC=Respective costs*Respective weight

=(0.75*11.25)+(0.05*5.5)+(0.2*3.965)

=9.51%(Approx).